INAF queue observing, 2016 July 3 - 4

Observer: F. Cusano, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O Kuhn on call
Instrument: LBC

Summary

At the sunset the sky is clear.
Still problem with the Dx AdSec we can not use LUCI2.
We use LBCs.

Night duration (12\xB0 el twilight) 20:30-4:00 ( 03:30-11.00 UT) = 7h 30m

Observing time: 6 h 10m (82 %)

Weather time loss: 0 h (0%)

technical time loss 0 h (0%)

note: 1h 20 m for LUCI+AO commisioning

Data Summary

Project_RUN Objects Filter ExpTime Notes
Momany NGC6205 SDT_U + r-sloan 3h some cirrus passing, seeing ~1
Mignani J2043+2740 SDT_U,V + V,i-sloan 1h  
LBC flats        

Details

2:40 starting sky flat with LBC, LBCB: SDT-U, V; LBCR: V, i

3:20 problem with LBC-R

3:30 problem solved restarting the LBC

3:36 slew to copointing of Momany

3:41 dofpia, X2, /first
  • seeing measured on the copoting images 0.8"

4:00 executing the firt ob of Momany, Science_NGC6205.ob
  • seeing 1" measured on science images at 4:05
  • seeing 1.1" measured on science images at 4:30

4:37 dofpia, X2

4:45 executing the second time Science_NGC6205.ob
  • seeing 1" measured on science images at 4:46
  • seeing 1.1" measured on science images at 4:55
  • seeing 0.9" measured on science images at 5:05

5:20 dofpia, X2
  • seeing measured on the copoting images 0.8"

5:25 executing the third time Science_NGC6205.ob
  • seeing 1.1" measured on science images at 5:26
  • seeing 1.2" measured on science images at 5:44

6:00 dofpia, X2

6:08 executing the 4th time Science_NGC6205.ob
  • some thin cirrus passing
  • seeing 0.9" measured on science images at 6:10

6:43 seeing stable <1.1".
  • We stop observing and give the telescope to AO commissioning

7:50 we can use again the telescope

8:00 slew to program of Momany, copointing ob
  • 8:01 dofpia

8:21 executing the 5th time Science_NGC6205.ob
  • seeing 1" at 8:22
  • seeing jump 1.5" at 8:32 due to passing cirrus
  • seeing again at 1" at 8:35
  • seeing 1" measured on science images at 8:51

* 8:56* dofpia

9:01 executing the 6th time Science_NGC6205.ob
  • seeing 1.1" at 9:05 on image
  • seeing jump 2.0" at 9:12 due to passing cirrus
  • seeing 1.1" at 9:18 on image

9:39 we slew to Mignani B3, copoiting
  • dofpia

9:54 sent 2015B_2_B3_Cal.ob, Mignani B3
  • script stopped condition not optimal for calibration

10:00 sent 2015B_2_B3_1.ob, Mignani B3
  • seeing 1.2"
  • seeing 0.7" at 10:08 on image
  • seeing 1.1" at 10:25 on image

10:45 sent dofpia and 2015B_2_B3_2.ob, Mignani B3
  • seeing 0.8" from dofpia
  • seeing 0.9" at 10:52
  • we stop the ob at 11:06 sky counts are too high

11:20 starting sky flats
Topic revision: r13 - 15 Dec 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback